OpenStudy (hba):

The amount of water in the tank is 't' after it has started to drain is given by w=100(t-15)^2 gallon.The water is running out after 5 min at rate of ?

OpenStudy (anonymous):

diffrentiaate w.r.t time

OpenStudy (hba):

w=100(t-15)^2 dw=200(t-15)dt

OpenStudy (hba):

so t=5 dw/dt=200(5-15) =-2000 gal/min

OpenStudy (anonymous):

so min is 5 u can solve @hba

OpenStudy (anonymous):

yup

OpenStudy (hba):

But the - sign shows that the water was running out ? so 2000 gal/min ?

OpenStudy (hba):

@Galaxyy

OpenStudy (anonymous):

exactly

OpenStudy (hba):

Thanks.